# Design criteria for zero-strain behavior in iron-based fluorides of tungsten bronze type as cathode materials for alkali ion batteries - a first-principles study

### v1

@@ II.2 Computational details
All calculations were spin-polarized (collinear). For all the compounds analyzed, comparative calculations showed that the iron ions are present in the high-spin state (i.e. maximizing the number of singly occupied 3d orbitals resulting in the highest possible magnetic moments per iron ion for the given number of electrons).
@@ III.1 Magnetic ordering of the Fe ions
$\Delta V_{0}=\frac{V_{0}^{\mathrm{final}}-V_{0}\mathrm{{}^{initial}}}{V_{0}\mathrm{{}^{initial}}}.$ (1)
@@ III.2 TTB NaxFeF3
We start our investigation with the TTB NaxFeF3 and compare the results with the experimental data of Han et al [16]. Since there are 20 possible intercalation sites in the chosen TTB supercell, 20! calculations would be necessary to determine the energetic order in which the individual sites would be occupied. Since this is impractical, we use the approach that first all sites of the site type with the lowest formation energy $E\mathrm{{}_{form}^{Na^{+}}}$ are filled, then the sites with the next lowest formation energy in the now partially occupied structure, and finally the sites of the remaining site type. The formation energy per Na cation is shown in Fig. 3(a) for the three intercalation steps. In the first step the 5-sites are occupied. The 4-sites have the lower formation energy in the 5-site occupied structure and are therefore occupied next, followed by the 3-sites.
Figure 3(b) shows the voltage profile for this stepwise occupation, w.r.t. Na/Na+, i.e. assuming a metallic sodium anode. The first (short) horizontal line corresponds to the occupation of one single 5-site and the last (short) horizontal line to all sites being occupied, except one single 3-site. The very small differences to the average values of the occupation of all resepective sites demonstrate the validity of the above described procedure of stepwise site-type occupation. The average voltage values are in good agreement with the values of Han et al. [16] who reported voltages between 2.5 and 4 V in the concentration range $x$=0 to $x$=0.6.
@@
The global volumes (represented by uppercase $V$’s) up to a concentration of $x$=0.6 are shown in Fig.  4(a). Global in this context means the volume of the supercell and thus stands for the expansion of a hypothetical perfect single crystal. The occupation of the 5-sites initially leads to a reduction in volume, while the subsequent occupation of the 4-sites leads to an expansion in volume. Overall, this results in a global volume change of 2.2% for a capacity comparable to that measured by Han et al. ($\sim$150 mAh/g). This result is in good agreement with the reported near-ZS behavior for this compound.
@@
During the intercalation of sodium into the sites, the volumes of the $A$Fj polyhedra change. These changes are displayed in Figure 4(b), where we show here and in the following the average volume per site type for better visibility (the typical variance between sites of the same site type is on the order of tenths of Å3). For both, 4-sites and 5-sites, the volume is smaller after occupation by Na than before. This can be explained by an electrostatic argument: There is an attraction between the positively charged Na+ cations and the negatively charged F- anions, and the intercalated sodium cation reduces the mutual repulsion of neighboring fluorine anions. This reduction in the local volume can partially compensate for the expansion of the iron octahedra, resulting in a small global volume expansion of the unit cell. For the occupation of the 5-sites this effect is more pronounced and therefore even leads to a global volume reduction, while the occupation of the 4-sites only results in a low local volume contraction and therefore does not completely compensate the expanding iron octahedra.
Figure 4: (a) Unit cell volumes for NaxFeF3 in the TTB structure. (b) Change in local volumes of the NaxFeF3 structure. For the FeF6 octahedra initial refers to the oxidation state Fe3+ and final to Fe2+. For the intercalation sites, inital and final refer to the average volume per site type of the F polyhedra surrounding the sites before and after occucpation by Na cations, respectively.
@@ IV.1 Suitability of TB structures as cathode materials
In the HTB structure, there is a compensation mechanism for the occupation of 6-sites, which leads to global volume changes of -1.5% and -0.2% for $A$=Li and $A$=Na, respectively. However, there is only one 6-site per unit cell, which corresponds to $x$=1/3 per formula unit. For the compound FeF3-0.33H2O, Li et al. reported a volume expansion of $\sim$3.9during the intercalation of $\Delta x$=0.66 Li and found indications that up to two Li+ cations occupied one 6-site [21]. We have not considered the possibility of double occupancy in this work, but we expect an expansion of the local $A$Fj volume (as opposed to contraction with single occupancy reported here) and thus a global expansion of the structure. Consequently, only low capacities with simultaneous ZS behavior are achievable in the HTB structure.
For TTB NaxFeF3 the ZS behavior was already measured experimentally and explained above by our atomistic simulations. In the case of K+, there is no site type that provides a sufficient compensation mechanism to maintain small global volume changes. For Li, however, both occupations of the 4-sites and of the 5-sites lead to a global volume reduction, while the occupation of the 3-sites expands the cell volume. This results in a small global volume change of 0.4% for the entire concentration range. However, if, as described here, the intercalation would strictly follow the site types, cell volume changes that go far beyond ZS behavior would be expected for the intermediate concentration steps. However, since this also applies to TTB NaxFeF3, where small changes were consistently measured in the experiments, it seems reasonable to expect a more mixed occupation and hence less pronounced global volume changes to a certain extent for TTB LixFeF3 as well. Temperature as well as kinetic effects may cause a more mixed occupation of sites. Thus, this compound represents a promising candidate material for ZS applications, with a theoretical capacity of 223 mAh/g and an average voltage of 3.23 V.
@@ IV.2 ZS design criteria
We used various features (e.g., bond lengths, symmetry and density features) as input for a number of structure descriptors which are implemented in the Python package matminer [46]. Those desciptors take the structural features as input and return a real number, which than can be used in further analysis. Each of these descriptors was then tested separately together with two parameters, the ionic radius of the $A$ cation $r\mathrm{{}_{ionic}^{\textit{A}}}$ and the unoccupied site volume $v\mathrm{{}_{unocc.}}$, in a linear regression model to find that structure descriptor, which results in the best prediction (i.e., lowest residual sum of squares, $R^{2}$) of the set of local volumes compared to the DFT data. Applying the Python package sklearn for this task [47], the descriptor MAD (mean absolute deviation of relative bond length) was identified as the optimal choice, resulting in a $R^{2}$ of 0.986. The linear equation of the predictor for standardized input data (removing the mean value and scaling to unit variance) was obtained as follows:
@@
Although the proposed descriptor has only been trained on iron-based fluorides of tungsten bronze type structures and thus the applicability to other structures and/or other transition metal-based fluorides/oxides is not granted, we are confident that it represents a step towards quantitative design criteria for ZS behavior. In combination with the systematic investigation of the influence of 3d TM on global volume changes in Ref. [15], a simple model is conceivable that can predict e.g. the global volume changes in other (i.e., non-iron-based) TB structures. Furthermore, the extension to other TM oxide/fluoride structures is well conceivable, possibly using some other descriptors, but still representing a computational inexpensive step compared to e.g., DFT calculations. However, this model is not expected to be a complete replacement for DFT calculations, since, e.g., the determination of the magnetic ground state still relies on a calculation of the electronic structure.

### v2

# First-principles study of strain behavior in iron-based fluorides of tungsten bronze type as cathode materials for alkali ion batteries

@@ II.2 Computational details
All calculations were spin-polarized (collinear). For all the compounds analyzed, comparative calculations showed that the iron ions are present in the high-spin state (i.e. maximizing the number of singly occupied 3d orbitals resulting in the highest possible magnetic moments per iron ion for the given number of electrons). The distribution of Fe2+ and Fe3+ ions was not specified prior to relaxation. After the relaxation of the ionic and electronic degrees of freedom, Fe ions were identified as Fe2+ or Fe3+ from their final magnetic moments.
@@ III.1 Magnetic ordering of the Fe ions
$\Delta V_{0}=\frac{V_{0}^{\mathrm{final}}-V_{0}\mathrm{{}^{initial}}}{V_{0}\mathrm{{}^{initial}}}\times 100\%.$ (1)
@@ III.2 TTB NaxFeF3
We start our investigation with the TTB NaxFeF3 and compare the results with the experimental data of Han et al [16]. The 20 possible intercalation sites in the TTB supercell would result in too many possible inequivalent configurations to be treated within the scope of this work to determine the energetic order of occupation of individual sites. Therefore, we use the approach that first all sites of the site type with the lowest formation energy $E\mathrm{{}_{form}^{Na^{+}}}$ are filled, then the sites with the next lowest formation energy in the now partially occupied structure, and finally the sites of the remaining site type. The formation energy per $A$ cation is defined as follows:
$E_{\mathrm{form}}^{A}=\frac{E_{\mathrm{final}}^{A_{x}\mathrm{FeF_{3}}}-E_{\mathrm{initial}}^{A_{x}\mathrm{FeF_{3}}}-nE_{\mathrm{bulk}}^{A}}{n}.$ (2)
Here, $E_{\mathrm{initial}}^{A_{x}\mathrm{FeF_{3}}}$ and $E_{\mathrm{final}}^{A_{x}\mathrm{FeF_{3}}}$ are the energies at lower and higher $x$ concentrations, respectively, and $n$ is the number of intercalated cations, related to $\Delta x$ by $\Delta n=x\times$formula units. The formation energy per Na cation is shown in Fig. 3 (a) for the three intercalation steps. In the first step the 5-sites are occupied. The 4-sites have the lower formation energy in the 5-site occupied structure and are therefore occupied next, followed by the 3-sites.
Figure 3 (b) shows the voltage profile for this stepwise occupation, w.r.t. Na/Na+, i.e. assuming a metallic sodium anode. The first (short) horizontal line corresponds to the occupation of one single 5-site and the last (short) horizontal line to all sites being occupied, except one single 3-site. The very small differences to the average values of the occupation of all resepective sites demonstrate the validity of the above described procedure of stepwise site-type occupation. The average voltage values are in good agreement with the values of Han et al. [16] who reported voltages between 2.5 and 4 V in the concentration range $x$=0 to $x$=0.6. In agreement with the geometric arguments given in that paper that the 3-sites are too small for suitable Na+ storage, the occupation of this site-type showed the lowest formation energy.
@@
The global volumes (represented by uppercase $V$’s) up to a concentration of $x$=0.6 are shown in Fig.  4 (a). Global in this context means the volume of the supercell and thus stands for the expansion of a hypothetical perfect single crystal. The occupation of the 5-sites initially leads to a reduction in volume, while the subsequent occupation of the 4-sites leads to an expansion in volume. Overall, this results in a global volume change of 2.2% for a capacity comparable to that measured by Han et al. ($\sim$150 mAh/g). This result is in good agreement with the reported near-ZS behavior for this compound. In reality the structure is rarely a perfect single crystal, and thermal and kinetic effects play a role. If the 4-sites were occupied first, the material would expand to a volume of 62.4 Å per unit cell, as indicated in Fig.  4 (a) with dotted lines. Occupying the 3-sites as first sites would lead to an even larger volume expansion of 67.6 Å per unit cell, but this is associated with a formation energy of -2.7 eV/atom as opposed to -3.3 eV/atom for the 5-sites. Assuming a Boltzmann distribution, at a temperature $T$ the probability of finding a system in a state with energy $E$ is proportional to $e^{(-E/k\mathrm{{}_{B}}T)}$. At room temperature $k\mathrm{{}_{B}}T$ is about 0.025 eV, which is an order of magnitude smaller than the differences in formation energies. Therefore, we will take the lowest states as a good first approximation to the behavior of the material, but note that in reality a more mixed occupation can be expected for intermediate concentrations, e.g., some occupation of 4-sites in TTB NaxFeF3 before all 5-sites are occupied, which would lead to a volume change curve lying between the two limit cases outlined in Fig.  4 (a).
@@
During the intercalation of sodium into the sites, the volumes of the $A$Fj polyhedra change. These changes are displayed in Figure 4 (b), where we show here and in the following the average volume per site type for better visibility (the typical variance between sites of the same site type is on the order of tenths of Å3). Unoccupied indicates the local volume of the site-type in the structure from the immediately preceding concentration, i.e. the global structure may be partially occupied. For both, 4-sites and 5-sites, the volume is smaller after occupation by Na than before. This can be explained by an electrostatic argument: There is an attraction between the positively charged Na+ cations and the negatively charged F- anions, and the intercalated sodium cation reduces the mutual repulsion of neighboring fluorine anions. This reduction in the local volume can partially compensate for the expansion of the iron octahedra, resulting in a small global volume expansion of the unit cell. For the occupation of the 5-sites this effect is more pronounced and therefore even leads to a global volume reduction, while the occupation of the 4-sites only results in a low local volume contraction and therefore does not completely compensate the expanding iron octahedra.
Figure 4: (a) Unit cell volumes for NaxFeF3 in the TTB structure. The dotted line shows the energetically less favorable course if the 4-sites were occupied first and then the 5-sites. (b) Change in local volumes of the NaxFeF3 structure. For the FeF6 octahedra initial refers to the oxidation state Fe3+ and final to Fe2+. For the intercalation sites, inital and final refer to the average volume per site type of the F polyhedra surrounding the sites before and after occucpation by Na cations, respectively.
@@ IV.1 Suitability of TB structures as cathode materials
In the HTB structure, there is a compensation mechanism for the occupation of 6-sites, which leads to global volume changes of -1.5 % and -0.2 % for $A$=Li and $A$=Na, respectively. However, there is only one 6-site per unit cell, which corresponds to $\Delta x$ =1/3 per formula unit. This means that there is a global volume increase below $x$=2/3 (occupation of the 3-sites) and the small volume changes occur above $x$=2/3. For the compound FeF3-0.33H2O, Li et al. reported a volume expansion of $\sim$3.9during the intercalation of $\Delta x$=0.66 Li and found indications that up to two Li+ cations occupied one 6-site [21]. We have not considered the possibility of double occupancy in this work, but we expect an expansion of the local $A$Fj volume (as opposed to contraction with single occupancy reported here) and thus a global expansion of the structure. Consequently, only low capacities with simultaneous ZS behavior are achievable in the HTB structure.
For TTB NaxFeF3 the ZS behavior was already measured experimentally and explained above by our atomistic simulations. In the case of K+, there is no site type that provides a sufficient compensation mechanism to maintain small global volume changes. For Li, however, both occupations of the 4-sites and of the 5-sites lead to a global volume reduction, while the occupation of the 3-sites expands the cell volume. This results in a small global volume change of 0.8 % for the entire concentration range. For intermediate concentration changes, we calculated larger volume changes, far beyond ZS behavior, when the occupation occurs strictly stepwise by site-type (e.g., $\Delta V_{0}$=-7.8from x=0 to x=0.4 for the occupation of the 5-sites, see Fig. 5 (c), left). However, as stated above, it is reasonable to expect a more mixed occupation, and hence less pronounced global volume changes at intermediate concentrations (see also Fig. 4 (a)). Thus, TTB LixFeF3 represents a promising candidate material for ZS applications, with a theoretical capacity of 223 mAh/g and an average voltage of 3.23 V. For smaller cations, which show global volume contractions upon intercalation of certain site-types, a ZS mechanism can be achieved due to mixed occupation and hence the iron-based fluorides TB structures are not suitable ZS materials for the large K+ cations and for Na+ cations only for limited concentration ranges.
@@ IV.2 ZS design criteria
We used various features (e.g., bond lengths, symmetry and density features) as input for a number of structure descriptors which are implemented in the Python package matminer [46]. Each descriptor takes a dataset of structural features as input and returns a numerical digit describing some aspect of the global structure. These values can then be used in numerical processing, such as analysis or building predictive models. Each of these descriptors was tested separately together with two parameters, the ionic radius of the $A$ cation $r\mathrm{{}_{ionic}^{\textit{A}}}$ and the unoccupied site volume $v\mathrm{{}_{unocc.}}$, in a linear regression model to find that structure descriptor, which results in the best prediction (i.e., lowest residual sum of squares, $R^{2}$) of the set of local volumes compared to the DFT data. Applying the Python package sklearn for this task [47], the descriptor MAD (mean absolute deviation of relative bond length) was identified as the optimal choice, resulting in a $R^{2}$ of 0.986. The linear equation of the predictor for standardized input data (removing the mean value and scaling to unit variance) was obtained as follows:
@@
Although the proposed descriptor was only trained on iron-based fluorides with TB-type structures and is therefore not directly applicable to other transition metal-based fluorides/oxides, it shows that simple correlations can be developed for such systems and help to find possible ZS materials. We are confident that it represents a step towards quantitative design criteria for ZS behavior. In combination with the correlation found in Ref. [15] between changes in the size of TM ions (Shannon radii) and the local volumes of the anion octahedra around them, a more complex but still computational inexpensive model can be envisioned that can predict possible ZS candidates in the subset of the 3d TM TB. Such models are not expected to be a complete replacement for DFT calculations, since, e.g., the determination of the magnetic ground state still relies on a calculation of the electronic structure but constitute a promising first screening step.